OpenStudy (anonymous):

Determine an equation of a circle that satisfies the given conditions. Answers must be in standard form. Center (2,8) passes through (-1,7)

OpenStudy (anonymous):

The equation of a circle is given by \[\left( x-h \right)^{2} + \left( y-k \right)^{2} = r ^{2}\] where (h,k) is (x,y) coordinates for the centre of the circle, and r is the radius of the circle. Using this information and the fact that the circle passes a point where x = -1 and y = 7, you can find the value of r. Once you have h, k and r, you have the equation of the circle.

OpenStudy (anonymous):

Also, when you're looking for the value of r, remember that it has to be positive because r is really a distance (the radius is the distance from the centre of the circle to its circumference) and because distances have to be positive, r would also have to be evaluated as positive.

OpenStudy (anonymous):

Can I use the distance formula to find r?

OpenStudy (anonymous):

You could, actually. I didn't even think of that! I was suggesting that you plug in the values of x = -1 and y = 7 into the circle equation (as well as the h and k values) so that you're left with an equation where the only variable is r^2, then solve for r. However, the distance formula should work too!

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ 7}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 8}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2} \\ \quad \\ (x-{\color{brown}{ h}})^2+(y-{\color{brown}{ h}})^2=r^2\qquad center\ ({\color{brown}{ h,k}})\quad radius=r\)

OpenStudy (jdoe0001):

hmm sorta forgot the "k" anyhow \(\bf (x-{\color{brown}{ h}})^2+(y-{\color{brown}{ k}})^2=r^2\qquad center\ ({\color{brown}{ h,k}})\quad radius=r\)

OpenStudy (anonymous):

for the distance formula is got 7.62 (I rounded it)

OpenStudy (anonymous):

@jdoe0001 and @Jesusc

OpenStudy (anonymous):

Can we see your working? Also, try not to round it - it's usually preferred to have the answer in exact form, i.e. leave it in surd form.

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ 7}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 8}}) \end{array}\qquad d = \sqrt{({\color{red}{ 2}}-{\color{red}{ (-1)}})^2 + ({\color{blue}{ 8}}-{\color{blue}{ 7}})^2}=\to ?\)

OpenStudy (jdoe0001):

well... that's not well written.. hamm

OpenStudy (jdoe0001):

actualy is ok anyhow is not 7.62 though

OpenStudy (jdoe0001):

notice that 8-7 = 1

OpenStudy (anonymous):

Lol I accidently wrote 8-1 instead of 8-7

OpenStudy (anonymous):

But I got 3.16227766

OpenStudy (anonymous):

Ok, cool. That's a correct approximation but try to leave it in surd form - the way I like to do this is by using my calculator (sometimes my brain) to work out what's under the surd ONLY. Then, I leave my answer with that number in the surd and that's it.

OpenStudy (anonymous):

So when I leave it in surd form that'll be r @Jesusc ?

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ 7}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 8}}) \end{array} \\ \quad \\ d = \sqrt{({\color{red}{ 2}}-{\color{red}{ (-1)}})^2 + ({\color{blue}{ 8}}-{\color{blue}{ 7}})^2}\to \sqrt{(3)^2+(1)^2}\to \sqrt{10} \\ \quad \\ \begin{array}{llll} (x-{\color{brown}{ 2}})^2+(y-{\color{brown}{ 8}})^2=10\\ \bf (x-{\color{brown}{ 2}})^2+(y-{\color{brown}{ 8}})^2=(\sqrt{10})^2 \end{array} \qquad center\ ({\color{brown}{ 2,8}})\quad radius=\sqrt{10}\)

OpenStudy (anonymous):

Yes, the surd form would be r (just as jdoe0001 has illustrated).

OpenStudy (anonymous):

Thank you guys for showing me instead of jut giving me the answer! I truly appreciate it. @jdoe0001 @Jesusc

OpenStudy (anonymous):

You're welcome :)

OpenStudy (jdoe0001):

yw